OpenStudy (anonymous):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle need an example to plzz

OpenStudy (anonymous):

if you know two side lengths of any right triangle you can always find the third. as far as the angles are concerned since you already know 1 of 'em (90 degrees) you if you can find one more, the other angle would be its supplement.

OpenStudy (callisto):

yes, if you know 2 sides of the triangle. use the trigo function to find it

OpenStudy (anonymous):

you can then use trig to find the angles after u have your side(s).

OpenStudy (callisto):

 tanΦ = 4/3 tanθ = 3/4